STUDY OF LARGE MOLECULAR CLUSTERS IN He DROPLETS

\underline{MIKHAIL N. SLIPCHENKO}, HIROMICHI HOSHINA, ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089, USA.

The rotationally resolved spectra of the ν_{3} vibrational mode of (CH_{4})_{n} (n = 1 − 2 × 10^{3}) clusters in He droplets have been measured. With increasing of the cluster size the effective rotational constant of the CH_{4} molecules in the cluster decreases and for n > 500 the rotational structure collapses. This behavior suggests gradual quenching of the internal rotation of the CH_{4} molecules upon increase of the cluster size.